# AnySR: Realizing Image Super-Resolution as Any-Scale, Any-Resource

### v1

@@ 4.1 Experiment Settings
Dataset. For training, current practices [38] use DF2K (including DIV2K [8] and Flicker2K [59]) as the training set, or pre-training on ImageNet1K [60] to enhance performance. For fairness, we start directly from the pre-trained models already applied with these techniques and focus solely on retraining with the DIV2K dataset.
For evaluation, we assess the PSNR on the most common evaluation datasets in the literature, including Set5 [53], Set14 [54], B100 [61], Urban100 [56], Manga109 [57].
@@
4.2 Main Results
Quantitative Results. Table 1 and Table 2 show the PSNR performances of off-the-shelf arbitrary-scale methods, compared to that of AnySR variants with different subnets dealing with individual scales and that of AnySR-retrained largest network for all scales. Fig. 4 presents a bar graph detailing the PSNR performances across all 30 scales ($\times$1.1$\sim$×$\times$4.0) for the three entities mentioned above. Considering the unavailability of some existing models (Meta-SR: EDSR; ArbSR: EDSR, RDN), we re-implement and re-train them following the experiment settings outlined in the papers [9, 51].
In Table 1, our AnySR approach markedly decreases computational expenses while preserving visual qualities, thereby improving network efficiency and scalability. In particular, when selecting the subnetwork for inference, there is an average performance drop of 0.15 dB on benchmark datasets, primarily observed at smaller scales. Considering the smaller subnetwork reduces computational costs by up to 50%, we find this performance drop acceptable. Additionally, Table 2 shows that, when using the entire network, the average performance drop is only 0.05 dB. This indicates that AnySR-retrained models can effectively preserve the original model’s performance during the rebuilding process, which can also be visually observed in Fig. 4. Such good performance is closely related to the reset probability $p$ in Line 4 of Algorithm 1, which will be further investigated in Sec. 4.4. Overall, as evidenced by the results, our AnySR succeeds in dynamically selecting the reconstruction network according to the availability of computing resources, thereby making better use of time-varying resources. Further details regarding network complexity will be explicitly discussed in Sec. 4.3.
@@ 4.2 Main Results
Figure 5: Visualization of existing arbitrary-scale SR models LIIF [10], SRNO [12], their AnySR variants (through different subnets) highlighted by , and AnySR-retrained version (through the largest network) denoted as .
@@
Qualitative Results. Fig. 5 compares the qualitative performances of LIIF [10], SRNO [12] and their corresponding AnySR variants on DIV2K [8] and Urban100 [56] under two different settings, i.e., different subnets to solve different scales and the entire network to solve all scales.
We show AnySR’s competitive visual quality against others, as there is little noticeable difference in the results. Notably, in the case of ×$\times 2$ scale, our method yields impressive visual qualities with substantially reduced computational overhead. From the visual results, it is noticeable that while AnySR exhibits slight deficiencies in saturation and spatial perception, the overall performance is commendable. The results with AnySR closely resemble the LIIF/SRNO image, featuring clear building edges and rich textural details on the leaves. Both AnySR and AnySR avoid generating artifacts or distortions. Therefore, our AnySR succeeds in rebuilding existing models to fulfill any-scale, any-resolution SR tasks while maintaining performance on par with the original models.
@@
Table 3: Comparison of the numbers of Parameters, PSNR performances, and FLOPs counts for SRNO [12] and its AnySR variants (through different subnets) highlighted by , AnySR-retrained version (through the largest network) denoted by . Parameters and FLOPs are measured under the setting of consistent pixel size for upsampling SR images on all scales.
@@
SRNO 1.50 (122.95%) 33.74 (--0.09) 97.93 (69.25%) 30.47 (--0.03) 55.54 (88.86%) 28.78 (--0.01) 35.06 (99.39%)
SRNO 1.50 (122.95%) 33.81 (--0.02) 140.40 (99.27%) 30.49 (--0.01) 62.12 (99.39%) 28.78 (--0.01) 35.06 (99.39%)
4.3 Complexity Analysis
@@
We perform a thorough analysis and comparison on SR tasks at different scales, encompassing the number of parameters, PSNR performance, and FLOPs, for SRNO [12] and its AnySR variants of different subnets to solve different scales and the entire network to solve all scales. All results are evaluated on top of the EDSR [1] backbone network on the Set14 dataset [54] as shown in Table 3.
In terms of parameters, our AnySR variant introduces only 0.28 M additional weights (22.95% of the original 1.22 M parameters), primarily dedicated to any-scale enhancement. In relation to computational costs, when performing the ×$\times 2$ task with chosen subnetworks, the total number of FLOPs (97.93 G), constitutes only 69.25% of the original 141.28 G. This results in a minimal performance loss of 0.09 dB, making it particularly suitable for devices with limited computational power. In the presence of ample computing resources, AnySR sustains an excellent performance-consumption balance on the original basis. Specifically, the results with AnySR in Table 3 across three scales reveal that the FLOPs make up only 99.27%, 99.39%, 99.39% of the original model, though accompanied by a negligible PSNR drop of 0.01$\sim$0.02 dB.
In summary, our AnySR is crafted as an inference-efficient method, facilitating efficient deployment on devices with time-varying resources while maintaining good performance, as demonstrated by extensive experiments.
4.4 Ablation Studies

### v2

@@ III-C Any-Scale Enhancement
[…]
MetaSR 0.021 0.051 0.075 0.037 0.082 0.120 0.053 0.098 0.132 0.037 0.088 0.130 0.037 0.032 0.053
MetaSR 0.022 0.052 0.074 0.038 0.085 0.121 0.055 0.100 0.132 0.038 0.091 0.129 0.038 0.033 0.053
MetaSR 0.021 0.050 0.074 0.037 0.083 0.121 0.053 0.098 0.132 0.037 0.088 0.129 0.037 0.032 0.053
LIIF 0.039 0.092 0.133 0.067 0.157 0.224 0.111 0.227 0.303 0.043 0.110 0.170 0.074 0.101 0.130
LIIF 0.040 0.093 0.136 0.071 0.165 0.233 0.113 0.233 0.312 0.046 0.122 0.187 0.074 0.103 0.135
LIIF 0.040 0.094 0.137 0.071 0.165 0.232 0.113 0.235 0.313 0.045 0.119 0.182 0.074 0.103 0.133
ArbSR 0.021 0.052 0.075 0.037 0.084 0.122 0.054 0.098 0.135 0.036 0.087 0.134 0.011 0.033 0.057
ArbSR 0.021 0.052 0.074 0.039 0.086 0.122 0.054 0.098 0.134 0.038 0.089 0.128 0.011 0.033 0.055
ArbSR 0.021 0.052 0.074 0.037 0.085 0.122 0.054 0.097 0.134 0.035 0.083 0.128 0.011 0.032 0.055
SRNO 0.039 0.090 0.133 0.066 0.157 0.222 0.105 0.222 0.297 0.039 0.106 0.162 0.015 0.043 0.073
SRNO 0.040 0.093 0.133 0.068 0.161 0.224 0.106 0.223 0.298 0.041 0.111 0.163 0.015 0.044 0.074
SRNO 0.039 0.091 0.133 0.067 0.160 0.224 0.106 0.224 0.298 0.040 0.107 0.163 0.015 0.043 0.074
IV Experimental Results
IV-A Experiment Settings
IV-A1 Training Details
We primarily rely on existing well-established arbitrary-scale SR models, including Meta-SR [9], LIIF [10], ArbSR [52], and SRNO [12], for AnySR rebuilding. Without loss of generality, experiments on top of two famous feature extraction backbones, including EDSR [1] and RDN [3], are conducted to validate the universality of AnySR.
During specific implementation, we configure the network number $T=4$ and $\{w_{i}\}^{T}_{i=1}=\{0.5,0.7,0.9,1.0\}$, that is, the smallest scale group $S_{1}$ enables a reduction in inference cost by up to 50% compared with the original. The reset probability $p$ in Line 4 of Algorithm 1 is configured at 0.6, and the hyper-parameter $\lambda$ in feature-interweaving is set as 4 for EDSR and 8 for RDN. Accordingly, the scale groups are designated as $\mathcal{S}_{1}=\{1.1,1.2,...,1.7\}$, $\mathcal{S}_{2}=\{1.8,1.9,...,2.5\}$, $\mathcal{S}_{3}=\{2.6,2.7,...,3.2\}$ and $\mathcal{S}_{4}=\{3.3,3.4,...,4.0\}$.
We train AnySR for 500 epochs from the pre-trained models, with an initial learning rate $10^{-5}$ decayed by 0.5 every 200 epochs. As pre-trained models for some existing methods (Meta-SR: EDSR; ArbSR: EDSR, RDN) are not available, we re-implement and re-train them following the experiment settings outlined in the papers [9, 52]. To ensure a fair performance evaluation, we keep identical settings and configurations used for the original models in our training process. The patch size is $50\times 50$ for Meta-SR and ArbSR, $48\times 48$ for LIIF, and $128\times 128$ for SRNO, with a batch size of 8 per GPU for EDSR and 4 per GPU for RDN. We employ an $\ell_{1}$ loss [1] and the Adam optimizer [59].
IV-A2 Dataset
For training, current practices [38] use DF2K (including DIV2K [8] and Flicker2K [60]) as the training set, or pre-training on ImageNet1K [51] to enhance performance. For fairness, we choose to follow the traditional training regimen directly on the DIV2K dataset to ensure a direct comparison with previous studies and to maintain consistency with the experimental protocols established in the literature.
For evaluation, we assess the PSNR [61] and LPIPS [62] on the most common evaluation datasets, including Set5 [54], Set14 [55], B100 [63], Urban100 [57], and Manga109 [58].
TABLE II: PSNR (dB) and LPIPS comparison for fractional-scale SR tasks of SRNO [12] and their AnySR variants, highlighted by (through different subnets) and (AnySR-retrained version via the largest network). AnySR performance is highlighted in blue if within 0.15dB (PSNR) / 0.01 (LPIPS) of the vanilla method, and red if improved.
Backbone Network Methods Set5 Set14 B100 Urban100 Manga109
$\times$1.8 $\times$2.6 $\times$3.6 $\times$1.9 $\times$2.9 $\times$3.8 $\times$1.9 $\times$2.7 $\times$3.5 $\times$1.8 $\times$2.6 $\times$3.7 $\times$1.3 $\times$2.6 $\times$3.3
[…]
SRNO 39.12 35.72 33.14 34.42 30.74 29.06 32.79 29.87 28.33 33.94 29.83 27.02 45.29 35.71 32.93
SRNO 39.04 35.68 33.12 34.34 30.70 29.04 32.73 29.84 28.31 33.60 29.72 26.99 45.01 35.57 32.87
SRNO 39.11 35.70 33.12 34.39 30.71 29.04 32.78 29.85 28.31 33.85 29.77 26.99 45.28 35.63 32.87
[…]
SRNO 0.028 0.070 0.119 0.048 0.127 0.203 0.078 0.184 0.274 0.031 0.089 0.154 0.010 0.032 0.065
[…]
Backbone Network Methods Set5 Set14 B100 Urban100 Manga109
$\times$1.8 $\times$2.6 $\times$3.1 $\times$1.8 $\times$2.7 $\times$3.5 $\times$1.8 $\times$2.6 $\times$3.7 $\times$1.8 $\times$3.1 $\times$3.8 $\times$1.6 $\times$1.8 $\times$3.4
[…]
SRNO 39.31 35.92 34.65 35.35 31.50 29.73 33.54 30.28 28.21 34.67 28.88 27.35 42.39 40.80 33.30
SRNO 39.21 35.86 34.51 35.24 31.40 29.73 33.46 30.20 28.20 34.31 28.63 27.32 42.20 40.61 33.27
SRNO 39.29 35.92 34.67 35.37 31.53 29.73 33.52 30.26 28.20 34.59 28.85 27.32 42.39 40.78 33.27
[…]
SRNO 0.028 0.068 0.115 0.047 0.123 0.198 0.075 0.178 0.268 0.026 0.080 0.141 0.010 0.030 0.062
[…]
@@ IV-A2 Dataset
TABLE III: PSNR (dB) and LPIPS comparison for out-of-distribution scales of arbitrary-scale SR methods and their AnySR variants, highlighted by (through different subnets) and (AnySR-retrained version). AnySR performance is marked in blue if within 0.15dB (PSNR) / 0.01 (LPIPS) of the baseline, and red if improved.
Set5 Set14 Urban100
Backbone Method $\times$6 $\times$12 $\times$18 $\times$24 $\times$30 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
[…]
MetaSR 28.48 24.17 22.01 20.97 20.06 26.30 22.99 21.44 20.43 19.60 23.57 20.74 19.47 18.71 18.17
[…]
LIIF 28.96 24.56 22.48 21.37 20.58 26.45 23.16 21.61 20.66 19.81 23.79 20.89 19.62 18.84 18.29
[…]
ArbSR 28.39 23.76 21.69 20.64 19.75 26.19 22.57 20.99 19.93 19.22 23.47 20.37 19.08 18.33 17.82
[…]
SRNO 29.06 24.62 22.52 21.37 20.54 26.55 23.19 21.65 20.68 19.77 24.08 21.10 19.75 18.96 18.40
[…]
MetaSR 0.123 0.211 0.263 0.281 0.297 0.194 0.266 0.302 0.319 0.333 0.224 0.340 0.383 0.400 0.410
[…]
LIIF 0.204 0.379 0.510 0.587 0.626 0.330 0.487 0.567 0.625 0.672 0.287 0.489 0.603 0.668 0.709
[…]
ArbSR 0.133 0.265 0.319 0.333 0.343 0.180 0.279 0.318 0.330 0.339 0.229 0.350 0.389 0.401 0.404
[…]
SRNO 0.210 0.394 0.525 0.619 0.636 0.332 0.496 0.579 0.637 0.687 0.281 0.491 0.610 0.677 0.719
[…]
MetaSR 28.95 24.36 22.04 20.99 20.01 26.57 23.10 21.50 20.51 19.55 24.06 20.96 19.62 18.83 18.25
[…]
LIIF 29.15 24.86 22.66 21.50 20.57 26.64 23.24 21.73 20.78 19.85 24.20 21.15 19.80 19.00 18.44
[…]
ArbSR 28.45 23.57 21.53 20.45 19.55 26.18 22.44 20.83 19.79 19.09 23.55 20.27 18.97 18.24 17.73
[…]
SRNO 29.40 24.94 22.73 21.50 20.65 26.76 23.33 21.82 20.78 19.96 24.42 21.35 19.95 19.15 18.55
[…]
MetaSR 0.114 0.195 0.248 0.272 0.295 0.184 0.257 0.293 0.308 0.326 0.199 0.320 0.370 0.389 0.403
[…]
LIIF 0.197 0.361 0.491 0.566 0.601 0.319 0.472 0.551 0.604 0.657 0.263 0.457 0.574 0.644 0.690
[…]
ArbSR 0.132 0.267 0.316 0.333 0.340 0.179 0.278 0.316 0.327 0.335 0.229 0.351 0.388 0.398 0.401
[…]
SRNO 0.206 0.387 0.514 0.591 0.628 0.323 0.485 0.565 0.623 0.674 0.267 0.474 0.595 0.665 0.711
[…]
IV-B Main Results
IV-B1 Quantitative Results
Extensive experiments are conducted using in-distribution and out-of-distribution scales, and we analyze the quantitative performance below to show the efficacy of our proposed AnySR.
In-Distribution Scales. Table I presents the PSNR and LPIPS performance on integer scales, showing the comparison between AnySR variants with different subnets dealing with individual scales and AnySR-retrained largest network for all scales. Additionally, Table II shows the quantitative results on fractional scales, with three fractional scales randomly selected for each dataset on the SRNO model [12]. More PSNR results of AnySR across all 30 scales ($\times$1.1–$\times$4.0) are visualized in Fig. 4, to provide a comprehensive overview of the results.
From Table I and Table II, there is a limited performance drop at smaller scales with an average decrease of 0.15dB (PSNR) / 0.05 (LPIPS). With only this minimal performance loss, we can transform existing methods into an any-scale, any-resource implementation by dynamically selecting the network based on available computational resources, and we believe this trade-off is worthwhile. When selecting the entire network, AnySR shows only a negligible performance drop across all scales, with an average PSNR difference of merely 0.05dB, and the average LPIPS difference is even less than 0.01. This indicates that AnySR-retrained models can effectively preserve the original model’s performance during the rebuilding process, which can also be visually observed in Fig. 4. Such good performance is closely related to the reset probability $p$ in Line 4 of Algorithm 1, investigated in Sec. IV-E.
Out-of-Distribution Scales. Table III presents the performance of out-of-distribution scales. For scales near the training range ($\times$6, $\times$12), AnySR shows overall metric improvements. However, as the scales deviate from the training range, there is a noticeable performance drop. Nevertheless, we observe across all scales, the AnySR-retrained version achieves better LPIPS scores, indicating that the images generated by AnySR maintain a high level of visual quality. Overall, these results demonstrate that our AnySR preserves, and in some cases surpasses, the reconstruction capability of the original network, allowing for stable, high-quality outputs across devices with varying computational resources through the dynamic selection of different reconstruction networks.
@@ IV-B1 Quantitative Results
Figure 5: Visualization of existing arbitrary-scale SR models LIIF [10], SRNO [12], Meta-SR [9], ArbSR [52], their AnySR variants (through different subnets) highlighted by , and AnySR-retrained version (through the largest network) denoted as .
Figure 6: Visualization comparison of SRNO [12], its AnySR variant (highlighted by ), and AnySR-retrained version (denoted by ) in representative scenarios. Line 1: dense and fine grains/textures; Line 2: animal fur textures; Line 3: dense and complex lines.
Figure 7: Visualization results for out-of-distribution scales of existing arbitrary-scale SR models and their AnySR-retrained version (denoted by ).
IV-B2 Qualitative Results.
A large number of qualitative experiments have been carried out, demonstrating that AnySR can generate competitive, high-fidelity visual images while reducing computational costs.
@@
In-Distribution Scales. Fig. 5 makes a comparison between off-the-shelf arbitrary-scale SR models and their corresponding AnySR variants on DIV2K [8] and Urban100 [57] under the two specific settings including: different subnets for different scales and the entire network for all scales.
Qualitative results of the models are satisfactory, as both AnySR and AnySR avoid generating artifacts or distortions. As shown in Line 2 and Line 3 of Fig. 5, both clear building edges and rich textural details on the leaves are retained.
We conduct evaluations of the model’s reconstruction ability across different scales using representative scenarios. In Line 1 of Fig. 6, we task our AnySR with reconstructing fine-grained, densely arranged objects to assess its handling of detail, yielding commendable results. We can clearly observe the frost on blueberries, the dew on leaves, and the texture on coral surfaces. In Line 2, AnySR successfully reconstructs complex animal patterns (zebra stripes) and realistic, delicate animal fur (feathers of the swan and the eagle). In Line 3, our model accurately reconstructs the intricate and intertwined lines of buildings, resulting in clear and organized outputs.
All reconstructed results demonstrate the robust reconstruction capability of AnySR. Even at the $\times$2 scale, where computational costs are reduced by nearly 30%, the model still achieves visual results on par with the original model. This reflects AnySR’s ability to deliver high-quality reconstructed images, even under varying computational resource conditions.
[…]
@@
TABLE IV: Comparison of the numbers of Parameters, PSNR performances, and FLOPs count for existing arbitrary-scale SR methods and their AnySR variants (through different subnets) highlighted by , AnySR-retrained version (through the largest network) denoted by . Parameters and FLOPs are measured with consistent pixel size for upsampling SR images across all scales.
@@
MetaSR 1.22 (100.00%) 33.57 ($+$0.00) 140.20 (100.00%) 30.29 ($+$0.00) 62.03 (100.00%) 28.52 ($+$0.00) 35.01 (100.00%)
MetaSR 1.50 (122.95%) 33.44 (--0.13) 97.73 (69.71%) 30.28 (--0.01) 55.45 (89.40%) 28.55 ($+$0.03) 35.01 (100.00%)
MetaSR 1.50 (122.95%) 33.58 ($+$0.01) 140.20 (100.00%) 30.32 ($+$0.03) 62.03 (100.00%) 28.55 ($+$0.03) 35.01 (100.00%)
LIIF 1.22 (100.00%) 33.66 ($+$0.00) 140.20 (100.00%) 30.34 ($+$0.00) 53.03 (100.00%) 28.62 ($+$0.00) 35.01 (100.00%)
LIIF 1.50 (122.95%) 33.53 (--0.13) 97.73 (69.71%) 30.32 (--0.02) 45.23 (85.30%) 28.61 (--0.01) 35.01 (100.00%)
LIIF 1.50 (122.95%) 33.62 (--0.04) 140.20 (100.00%) 30.35 ($+$0.01) 53.02 (100.00%) 28.61 (--0.01) 35.01 (100.00%)
ArbSR 1.22 (100.00%) 33.70 ($+$0.00) 140.92 (100.00%) 30.31 ($+$0.00) 62.34 (100.00%) 28.56 ($+$0.00) 35.14 (100.00%)
ArbSR 1.50 (122.95%) 33.46 (--0.24) 101.86 (72.28%) 30.28 (--0.03) 57.26 (91.85%) 28.53 (--0.03) 35.14 (100.00%)
ArbSR 1.50 (122.95%) 33.65 (--0.05) 140.92 (100.00%) 30.30 (--0.01) 62.34 (100.00%) 28.53 (--0.03) 35.14 (100.00%)
@@
SRNO 1.50 (122.95%) 33.74 (--0.09) 97.93 (69.25%) 30.47 (--0.03) 55.54 (88.86%) 28.78 (--0.01) 35.28 (100.00%)
SRNO 1.50 (122.95%) 33.81 (--0.02) 141.28 (100.00%) 30.49 (--0.01) 62.50 (100.00%) 28.78 (--0.01) 35.28 (100.00%)
TABLE V: Comparison of the number of parameters and FLOPs of a single ResNet Block of EDSR [1] in SRNO [12], its AnySR variants and its AnySR-retrained version.
Setting Metric Conv Relu GAP FC ALL
$\times$2, SRNO Params (K) 73.86 0.00 - - 73.86
FLOPs (M) 8860.00 3.84 - - 8863.84
[…]
FLOPs (M) 6090.00 2.64 2.64 0.024 6095.30
[…]
FLOPs (M) 8860.00 3.84 3.84 0.035 8867.72
$\times$3, SRNO Params (K) 73.86 0.00 - - 73.86
FLOPs (M) 3920.00 1.70 - - 3921.70
[…]
FLOPs (M) 3500.00 1.51 1.51 0.031 3503.05
[…]
FLOPs (M) 3920.00 1.70 1.70 0.035 3923.44
$\times$4, SRNO Params (K) 73.86 0.00 - - 73.86
FLOPs (M) 2220.00 0.96 - - 2220.96
[…]
FLOPs (M) 2220.00 0.96 0.96 0.035 2221.96
[…]
FLOPs (M) 2220.00 0.96 0.96 0.035 2221.96
IV-C Complexity Analysis
@@
We perform a thorough comparison on SR tasks at different scales, encompassing the number of parameters, PSNR performance, and FLOPs, for existing arbitrary-scale SR methods and their AnySR variants of different subnets to solve different scales and the entire network to solve all scales. All results are evaluated on top of the EDSR [1] backbone network on the Set14 dataset [55] as shown in Table IV. Since the encoder contains 16 ResNet blocks, we calculate the parameters and FLOPs for a single image, for a single ResNet block under different settings, as shown in Table V. Taking SRNO as an example, we conduct a thorough and detailed analysis of the parameter count and computational cost.
In terms of parameters, our AnySR variant introduces only 0.28M additional weights (22.95% of the original 1.22M parameters). As shown in Table V, these additional weights come from the fully connected layers, which are primarily dedicated to any-scale enhancement.
As for computational cost, as shown in Table V, despite the introduction of fully connected layers and GAP operations, the FLOPs are almost negligible compared to those required by the convolutional layers of the original model. Instead, the selection of smaller subnets leads to significant savings in FLOPs. For example, when performing the ×$\times 2$ task with selected subnetworks, each ResNet block saves around 2800M FLOPs in the convolutional layers, while FLOPs introduced by the GAP operation and fully connected layers are less than 3M. The total number of FLOPs (97.93G) constitutes only 69.25% of the original 141.28G, with a minimal performance loss of 0.09 dB, making it suitable for devices with limited computational power to output high-quality results.
In case of ample computing resources, AnySR maintains a good performance-to-consumption balance. When performing the ×$\times 4$ task with the entire network, Table V shows that each ResNet block introduces negligible 0.995M FLOPs. For the entire backbone, the FLOPs of the AnySR-retrained version (35.28G) are approximately equal to those of the original model, implying no computing introduction.
TABLE VI: PSNR (dB) and LPIPS comparison of SRNO using smaller subnets (denoted as $F_{i}$) to perform out-of-distribution scale SR tasks. Underlines indicates the best performance.
Set5 Set14 Urban100
Backbone Method $\times$6 $\times$12 $\times$18 $\times$24 $\times$30 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
[…]
SRNO 29.06 24.62 22.52 21.37 20.54 26.55 23.19 21.65 20.68 19.77 24.08 21.10 19.75 18.96 18.40
$F_{1}$ 28.64 24.32 22.31 21.09 20.18 26.26 22.98 21.50 20.51 19.63 23.53 20.74 19.47 18.73 18.20
$F_{2}$ 28.92 24.55 22.44 21.26 20.41 26.46 23.10 21.58 20.62 19.72 23.83 20.92 19.61 18.82 18.28
$F_{3}$ 29.00 24.57 22.45 21.29 20.51 26.52 23.14 21.61 20.66 19.74 23.96 20.99 19.66 18.87 18.33
[…]
SRNO 0.210 0.394 0.525 0.619 0.636 0.332 0.496 0.579 0.637 0.687 0.281 0.491 0.610 0.677 0.719
$F_{1}$ 0.220 0.403 0.548 0.625 0.641 0.342 0.508 0.595 0.652 0.698 0.307 0.523 0.641 0.701 0.737
$F_{2}$ 0.213 0.398 0.526 0.624 0.650 0.333 0.498 0.581 0.640 0.689 0.289 0.497 0.617 0.683 0.725
$F_{3}$ 0.209 0.394 0.527 0.618 0.641 0.331 0.494 0.579 0.637 0.686 0.282 0.489 0.608 0.676 0.720
[…]
SRNO 29.40 24.94 22.73 21.50 20.65 26.76 23.33 21.82 20.78 19.96 24.42 21.35 19.95 19.15 18.55
$F_{1}$ 29.01 24.72 22.46 21.32 20.49 26.53 23.20 21.70 20.64 19.81 24.02 21.06 19.73 18.94 18.37
$F_{2}$ 29.20 24.81 22.56 21.28 20.56 26.65 23.24 21.72 20.68 19.82 24.22 21.18 19.80 19.00 18.41
$F_{3}$ 29.23 24.76 22.64 21.30 20.61 26.67 23.23 21.76 20.75 19.83 24.25 21.19 19.82 19.00 18.42
[…]
SRNO 0.206 0.387 0.514 0.591 0.628 0.323 0.485 0.565 0.623 0.674 0.267 0.474 0.595 0.665 0.711
$F_{1}$ 0.218 0.402 0.532 0.599 0.636 0.337 0.502 0.585 0.643 0.686 0.294 0.511 0.626 0.688 0.727
$F_{2}$ 0.210 0.375 0.501 0.571 0.614 0.328 0.484 0.562 0.619 0.666 0.274 0.476 0.590 0.658 0.704
$F_{3}$ 0.210 0.384 0.513 0.578 0.621 0.328 0.487 0.562 0.617 0.668 0.273 0.477 0.592 0.658 0.704
[…]
Figure 8: Visualization results of SRNO using smaller subnets (denoted as $F_{i}$) to perform out-of-distribution scale SR tasks on the EDSR backbone.
IV-D Subnet Exploration
We evaluate smaller subnets in handling high magnification factors. Table VI and Fig. 8 present the quantitative and qualitative results of SRNO and its AnySR variants with subnets of varying sizes, denoted as $F_{i}$, on out-of-distribution scales. Smaller subnets perform lower-scored PSNR, as reflected in the visual results with less detail recovery such as distant lighthouses and the fine edges of plant leaves. However, it improves as subnet size increases. The LPIPS scores show slight variation, indicating that AnySR’s subnets generally maintain good perceptual quality in the generated images.
IV-E Ablation Studies